# FLASC: A Flare-Sensitive Clustering Algorithm

### v1

@@ 6. Conclusion
We presented the FLASC algorithm that combines HDBSCAN* clustering with a graph-based branch-detection approach. We have shown that the algorithm can detect flare-based subgroups that do not contain local density maxima in real-world data without specifying features of interest or extracting the flares from a visualisation manually. In addition, we demonstrated that the branching hierarchy, as found by FLASC, can provide information about a cluster’s shape that is not present in HDBSCAN*’s output when the subgroups are detectable as clusters. Two analyses of synthetic data showed that FLASC provides similar results for samples of the same underlying distribution and that its computational performance scales similarly to HDBSCAN*.
Acknowledgements.
We thank Kris Luyten for his insightful comments on an early version of the manuscript. This work was supported by Hasselt University BOF grants [BOF20OWB33] and [BOF21DOC19].
Appendix A HDBSCAN* on branch stability data
Section 4.3.1 briefly mentioned that the data sets used for the stability analysis were designed so that the branches do not contain local density maxima. This section shows which subgroups HDBSCAN* detects on those data sets. HDBSCAN* was tuned with min samples $k=5$, min cluster size $m_{c}=8$, and the leaf cluster selection strategy. The parameters were selected using the procedure described in Section 4.3.1. Figure 11 shows the resulting classifications. Notice that subgroups are found within the branches, but the branches are not described by a single subgroup. In addition, many points are classified as noise by HDBSCAN*.
Figure 11. The data sets with varying noise ratios and branch lengths from Section 4.3 coloured by HDBSCAN*’s classification when tuned to detect multiple clusters. Branch lengths (columns) are normalised visually.
This figure shows the tree-pointed star point clouds for all branch length and noise ratio combinations with HDBSCAN*’s labelling indicating that it does not detect the branches as a single subgroup.
Figure 11. The data sets with varying noise ratios and branch lengths from Section 4.3 coloured by HDBSCAN*’s classification when tuned to detect multiple clusters. Branch lengths (columns) are normalised visually.
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" alt="[LOGO]">

### v2

@@ Conclusion
We presented the FLASC algorithm that combines HDBSCAN* clustering with branch-detection post-processing step. We have shown that the algorithm can detect branch-based subgroups that do not contain local density maxima in real-world data without specifying features of interest or extracting the branches from a visualisation manually. In addition, we demonstrated that branching hierarchies found by FLASC can provide information about the data’s shape that is not present in HDBSCAN*’s cluster hierarchy. Two synthetic benchmarks demonstrated FLASC’s stability and indicated FLASC’s computational performance scales similarly to HDBSCAN*.
Acknowledgements
We thank Kris Luyten for his comments on an early version of the manuscript.
Data Availability
The data and code used in this paper are available at https://zenodo.org/records/13326252.
Funding Statement
The work of D.M. Bot was supported by Hasselt University BOF grant [BOF20OWB33]. The work of J. Peeters was supported by Hasselt University BOF grant [BOF21DOC19]. The work of J. Aerts was supported by KU Leuven grant STG/23/040.